OpenStudy (anonymous):

Plzz help me How would you write the following expression as a sum or difference? log(3√2-x/3x) A.1/3(log(2-x)+log3x) B. 1/3log(2-x)-log3x C. 1/3log(2-x)-logx D. 1/3log(2-x)+log3x

OpenStudy (anonymous):

this is a review for a test

OpenStudy (anonymous):

\[\log\left(\frac{\sqrt[3]{2-x}}{3x}\right)\]?

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

or \[\log\left(\sqrt[3]{\frac{2-x}{3x}}\right)\]

OpenStudy (anonymous):

how did you write it like that

OpenStudy (anonymous):

magic

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

then start with \[\log(\sqrt[3]{2-x})-\log(3x)\]

OpenStudy (anonymous):

then go to \[\frac{1}{3}\log(2-x)-\log(3x)\]

OpenStudy (anonymous):

Ahh i see

OpenStudy (anonymous):

do u mind helping me with one more question?

OpenStudy (anonymous):

sure and btw i am writing it in \[\LaTeX\]